the withd of a rectangle is 3cm more than half the length. the perimeter is 96cm. what is the dimensions of the rectangle?
Answer by checkley77(12844)

W=L/2+3
2W+2L=96
2(L/2+3)+2L=96
L+6+2L=96
3L=96-6
3L=90
L=90/3
L=30 ANS. FOR THE LENGTH.
W=30/2+3
W=15+3=18 ANS. FOR THE WIDTH.
PROOF:
2*18+2*30=96
36+60=96
96=96